### v2

# Improving Time Series Forecasting via Instance-aware Post-hoc Revision

Abstract
Time series forecasting plays a vital role in various real-world applications and has attracted significant attention in recent decades. While recent methods have achieved remarkable accuracy by incorporating advanced inductive biases and training strategies, we observe that instance-level variations remain a significant challenge. These variations—stemming from distribution shifts, missing data, and long-tail patterns—often lead to suboptimal forecasts for specific instances, even when overall performance appears strong. To address this issue, we propose a model-agnostic framework, PIR, designed to enhance forecasting performance through Post-forecasting Identification and Revision. Specifically, PIR first identifies biased forecasting instances by estimating their accuracy. Based on this, the framework revises the forecasts using contextual information, including covariates and historical time series, from both local and global perspectives in a post-processing fashion. Extensive experiments on real-world datasets with mainstream forecasting models demonstrate that PIR effectively mitigates instance-level errors and significantly improves forecasting reliability. Our code is available [^1] https://github.com/icantnamemyself/PIR
1 Introduction
Time series forecasting is a fundamental task in time series analysis, attracting considerable attention in recent years [qiu2024tfb, wang2024deep]. Various applications have been facilitated by the advancement of forecasting, including traffic planning [jin2023spatio], stock market prediction [li2024master], healthcare analytics [kaushik2020ai], and weather forecasting [nature, nature2]. Recent years have witnessed significant efforts dedicated to this area, with deep learning-based approaches achieving remarkable success due to their powerful ability to capture both temporal [Informer, fedformer] and cross-channel dependencies [crossformer, itransformer]. Furthermore, advanced inductive biases and training strategies have been introduced to address the non-stationary nature of time series data [revin, san] and to construct foundation models for time series forecasting [timesfm, timer, moirai].
Despite their satisfactory performance in overall evaluations, we emphasize that existing forecasting approaches often overlook inherent instance-level variations, which can arise from the long-tail distribution of numerical patterns in time series data and ultimately lead to forecasting failures in specific cases. Specifically, time series typically represent the numerical reflections of complex and dynamic real-world systems and are therefore prone to noise, sensor failures, and other anomalies during data collection. These issues can result in potential distribution shifts [liu2024timeseries], missing values [tashiro2021csdi], or unforeseen anomalies [blazquez2021review]. Therefore, while most instances exhibit similar numerical patterns, a subset may present rare behaviors, and mainstream forecasting methods often struggle to effectively model these exceptional instances, leading to inaccurate or even unreliable forecasting outcomes.
Figure 1: Per-instance MSE evaluation of PatchTST on the ETTh1 dataset and its corresponding error distribution. The error varies among instances and exhibits a long-tail distribution due to the instance-level variations.
To better illustrate this phenomenon, we present the per-instance Mean Squared Error (MSE) evaluation curve of PatchTST [patchtst] on the ETTh1 dataset [Informer], along with the corresponding error distribution visualized through both a histogram and a kernel density estimate in Figure 1. As shown in the figure, while the MSE remains consistently low for the majority of instances, there exist specific cases where PatchTST yields unsatisfactory forecasting results, as indicated by multiple spikes in the error curve. Moreover, the error distribution clearly exhibits a long-tail pattern, reinforcing our motivation and underscoring the challenges posed by the instance-level variations.
To this end, we propose the PIR framework, designed to enhance forecasting results via Post-forecasting Identification and Revision, addressing the challenge from a novel post-processing perspective. The framework comprises two key components. The first is the failure identification mechanism, which identifies the potential biased forecasting instances where the model’s predictions are less reliable, through estimating the forecasting performance on a per-instance level. The second is a post-revision module, which refines the forecasts by leveraging contextual information from both local and global perspectives. For the local revision, inspired by exogenous variable modeling approaches [nbeatsx, timexer], we utilize immediate forecasts of covariates along with available exogenous information such as timestamps [wang2024rethinking] and textual descriptions [timemmd] as side information to implicitly mitigate the impact of instance-level variations within a local window. This approach is grounded in the assumption that dependencies between covariates like lead-lag effects can provide valuable insights into future trends [zhao2024rethinking], and exogenous information can serve as additional prior conditions guiding the revision. For the global revision, the framework addresses rare or atypical numerical patterns by explicitly retrieving similar instances from the global historical data [liu2024retrievalaugmented]. This retrieval-based strategy enables the model to better capture long-tail patterns that may be overlooked by conventional forecasting models. Finally, PIR integrates the original forecasts with its revision outputs via a weighted sum, making it model-agnostic and broadly compatible with existing forecasting architectures. The primary contributions of our work are summarized as follows:
  • We are the first to highlight the existence of instance-level variations that can lead to forecasting failures on certain cases, evident by the unsatisfactory performance of existing forecasting methods at the per-instance level.
  • We introduce PIR, a model-agnostic framework designed to address this challenge. The framework estimates the forecasting performance to identify the potential failures and utilizes contextual information to revise them from both local and global perspectives.
  • We conduct extensive experiments on well-established real-world datasets, covering both long-term and short-term forecasting settings with mainstream models. The results demonstrate that the PIR framework consistently enhances forecasting accuracy, leading to more reliable and robust performance.
2 Related Work
2.1 Time Series Forecasting
Time series forecasting has been a central area of research for several decades. One of the earliest landmark contributions was the development of statistical methods, which are celebrated for their solid theoretical foundation and systematic approach to model design. Representative works in this domain include ARIMA and Holt-Winters [arima, holt-winters], which have laid the groundwork for more advanced methodologies. More recently, with the advancement of deep learning, numerous neural forecasting models have been proposed, achieving superior performance. Recurrent Neural Networks (RNNs) are among the first deep learning architectures applied to forecasting [wen2017multi, deepar], followed by a variety of other structures, such as convolution-based models [lstnet, moderntcn, convtimenet], attention-based networks [logtrans, tft, fedformer], and MLP-based architectures [nbeats, DLinear, sparsetsf]. These models have demonstrated remarkable capabilities in capturing both temporal and cross-channel dependencies within time series data, leading to substantial improvements in forecasting accuracy.
Beyond advancements in model architecture, a range of specialized techniques rooted in time series analysis has also emerged. These include methods for time series decomposition [wu2021autoformer, yu2024revitalizing], frequency modeling [yi2024frequency], and approaches for non-stationary forecasting [revin, san], which address the unique challenges presented by time series data. Besides, self-supervised pretraining has gained significant attention as a powerful approach, with various training strategies being explored [woo2022cost, timemae, tpami_ssl]. In addition, recent efforts aim to construct foundational time series models capable of learning universal representations and being applied to diverse downstream tasks [gpht, timesfm, timer, moirai]. Moreover, some pioneer works explore forecasting enhanced with the reasoning ability of LLMs [wang2025can, luo2025time].
2.2 Context Modeling in Time Series Forecasting
In addition to forecasting based solely on multivariate time series data, several pioneering studies have explored the incorporation of contextual information to enhance forecasting performance [ashokcontext]. On one hand, some advancements focus on integrating exogenous information within a local window during the forecasting process. For instance, TFT [tft] and TiDE [tide] leverage dense encoders to process timestamp information, which is subsequently used to condition future forecastings. Similarly, NBEATSx [nbeatsx] and TimeXer [timexer] improves the forecasting accuracy of target variables by explicitly modeling the influence of exogenous variables. On the other hand, some studies investigate the potential of retrieving relevant time series from the global historical context. RATD [liu2024retrievalaugmented] utilizes the retrieved series as references to guide the denoising process of diffusion-based forecasters, and RATSF [wang2024ratsf] introduces the retrieval augmented cross-attention architecture for explicitly modeling similar historical data. More recently, RAFT [raft] achieves satisfactory performance through forecasting enhanced with multi-period retrieval.
Different from existing methods, our proposed PIR framework tackles a novel research challenge: mitigating the impact of instance-level variations that can lead to forecasting failures in certain cases. The framework begins by identifying the potential failure instances through estimating their accuracy, and then revises the forecasts by leveraging available contextual information from both local and global perspectives in a post-processing manner. As a result, PIR functions as a model-agnostic plugin, allowing it to be seamlessly integrated into arbitrary forecasting models for enhanced performance.
3 Methodology
In this section, we will delve into the specifics of the proposed PIR framework, which is illustrated in Figure 2. We begin with the problem definition, followed by a comprehensive description of the framework’s key components.
Figure 2: Overview of the proposed PIR framework. The framework first identifies the potential forecasting failure cases through estimating the performance of the intermediate results generated by backbone models on a per-instance level. Besides, the framework incorporates Local Revising and Global Revising components, which utilize contextual information, including available exogenous variables within the local window and global historical time series, to revise the forecasting results for enhanced performance.
3.1 Problem Definition
We first consider the general multivariate time series forecasting task. Given a set of input series $X=\{x_{i}\}_{i=1}^{M}$, the objective is to learn a mapping function $Y=f_{\theta}(X)$ that accurately forecasts the future values $Y=\{y_{i}\}_{i=1}^{M}$, where $x_{i}\in\mathbb{R}^{N\times L_{in}}$ represents the $i$-th input time series and $y_{i}\in\mathbb{R}^{N\times L_{out}}$ represents the corresponding target series. Here, $N$ denotes the number of channels and $M$ is the number of instances, while $L_{in}$ and $L_{out}$ denote the lengths of the input and target series, respectively. For simplicity and brevity, we will omit the indices for the time series instances in the following sections.
Moreover, the PIR framework focuses on a novel problem to revise the forecasting results for better performance. Specifically, given the intermediate results $\bar{Y}$ of arbitrary forecasting models, the objective is to learn the revising function $Y=f_{\phi}(X,\bar{Y},C)$ that corrects the potential forecasting failures. Here, $C$ represents the available contextual information, such as exogenous variables, that can aid in the revision process for more accurate and reliable forecasts.
3.2 Failure Identification
The goal of the PIR framework is to post-process forecasting results by identifying and revising potential forecasting failures caused by instance-level variations. Thus, the identification process becomes a primary target and challenge. Conceptually, it forms an Uncertainty Estimation task based on the immediate forecasting results. Though various efforts have been devoted in this area, it has not been deeply explored for properly estimating the uncertainty of point-wise forecasting results mainly due to two reasons. Firstly, given the regression nature of the forecasting task, most existing approaches generate predictions directly from the hidden states. As a result, token-level distributions before generation are not available, and common uncertainty evaluation methods based on probabilities are not directly applicable [perplexity, gal2016uncertainty]. Secondly, the forecasting failures arise from both data uncertainty (e.g., missing values) and model uncertainty (e.g., underfitting on long-tail patterns) [gawlikowski2023survey] while most existing uncertainty quantification methods primarily focus on the latter one.
Although we have identified several potential reasons that may lead to prediction failure, there are many other potential and interrelated factors, and there is a lack of ground truth and metrics to identify and disentangle the specific reasons for each forecasting failure instance. Therefore, we do not explicitly locate these failure patterns. Instead, we innovatively quantify uncertainty using forecasting error and adopt a data-driven approach as a practical solution for estimating the uncertainty of a given input-forecast pair $(x,\bar{y})$, which is more flexible and can potentially accommodate a broader range of failure modes. Specifically, following common practice in modeling complex dynamics [nst, fan2023dish], we employ a two-layer fully connected neural network with non-linear activation functions $f_{ue}(\cdot)$ to estimate the forecasting uncertainty. Additionally, we introduce an auxiliary constraint to support the abovementioned procedure. Since explicit uncertainty is not directly available for the forecasting results, we treat the forecasting error as a feasible guiding signal. The intuition behind this is that higher uncertainty will likely result in greater forecasting error. Therefore, the estimated uncertainty is further constrained to predict the MSE of the forecasting results:
$\delta$ $=f_{ue}(x,\bar{y},E),$ (1)
$\mathcal{L}_{ue}$ $=\frac{1}{N}\sum_{1}^{N}\|\delta-\|\bar{y}-y\|_{2}^{2}\|_{1}.$
Here $\delta$ is the estimated uncertainty, and $\|\cdot\|_{1}$ represents the Mean Absolute Error (MAE) loss function, which ensures that the estimated uncertainty aligns with the actual forecasting error. To provide additional contextual information, we introduce the channel embedding matrix $E=(e_{1},e_{2},...,e_{N})\in\mathbb{R}^{N\times d}$, which encodes the channel identity information. This matrix enables the model to better capture variations across different channels and provides crucial context for more accurate uncertainty estimation. Through this approach, the framework can generate uncertainty estimates specific to each model and input instance, thereby satisfying both types of uncertainty. Moreover, it forms a coupled framework, enabling efficient end-to-end optimization jointly with the forecasting task, and offers a degree of interpretability by using the forecasting error as a measurement for uncertainty.
3.3 Local Revising
After identifying the potential forecasting failures with high uncertainty, the PIR framework revises these results from both local and global perspectives. For the local revision component, the core idea is to leverage the contextual information within a local horizon window, which includes the intermediate forecastings of covariates and available exogenous variables, to enhance forecasting accuracy. The intuition behind this approach is straightforward. Firstly, the dependencies including lead-lag effects are widely present in the time series data, where the forecasting results of covariates can provide valuable insights into future trends [zhao2024rethinking]. Besides, the exogenous variables, such as the time-related information and environmental factors known in advance can serve as a prior condition [tft, wang2024rethinking], mitigating the impact of sudden distribution shifts caused by natural rhythms like holiday effects [harvey1990estimation, taylor2018forecasting]. Conceptually, this approach is particularly beneficial for models that prioritize robustness over capacity by adopting a channel-independent strategy [han2024capacity].
In practice, we project the intermediate forecasting results into hidden states on a per-variate basis [itransformer], along with the available exogenous variable information, which are concatenated for further correlation extraction. Let $h_{co},h_{exo}$ be the representation of covariates and exogenous variables respectively, and $c$ refer to the available exogenous variable information corresponding to the instance, the projection process is formulated as:
$H_{0}$ $=[h_{co},h_{exo}],$ (2)
$h_{co}$ $=\text{CoVariateEmb}(\bar{y}),$
$h_{exo}$ $=\text{ExoVariateEmb}(c).$
Here $\text{CoVariateEmb}(\cdot)$ is a trainable linear projector, and $\text{ExoVariateEmb}(\cdot)$ is implemented flexibly based on the characteristic of $c$, which can be a linear projector for numerical features or language models for textual descriptions [timemmd]. Through this approach, local context from multiple sources and domains can be seamlessly embedded in the framework to guide the revising process.
The hidden states are then passed through a traditional Transformer [vaswani2017attention] with a linear prediction head to generate the revised results $y_{local}$. By leveraging the attention mechanism, the local revision component explicitly captures the correlations between covariates and exogenous variables, thereby ensuring that the contextual information within the local window is fully utilized to correct forecasting failures and enhance prediction accuracy.
3.4 Global Revising
In addition, the PIR framework incorporates a global revision component that leverages global historical information to further refine forecasting results. As illustrated in Figure 1, instance-level variances can lead to a long-tail distribution in performance, primarily because traditional models struggle to capture rare numerical patterns. To address this challenge, we introduce a straightforward yet effective retrieval module that explicitly retrieves relevant historical time series exhibiting similar numerical characteristics. These retrieved series serve as reference signals during the revision process, enabling the model to better handle atypical patterns that are typically underrepresented in training.
Specifically, we first construct the retrieval database using only the training input-target time series pairs $(X_{train},Y_{train})$, as we treat historical information as the global context. This design not only prevents data leakage but also facilitates the extension of the database to incorporate multi-source datasets, since it depends solely on raw time series data. Based on the retrieval database, the top-$K$ most relevant time series are selected for a given input series $x$ by computing the similarity between $x$ and each candidate in the database, as formalized below:
$\text{Index},w=$ $\text{TopK }\operatorname{Sim}(Enc(x),Enc(X)),$ (3)
$Y_{re}$ $=\{Y_{train,i}|\forall i\in\text{Index}\}.$
Here, $Enc(\cdot)$ is an encoding function that processes the time series, which is instantiated as an instance normalization operation [revin] in practice for its simplicity and effectiveness of mitigating the impact of non-stationarity that could lead to unstable similarity estimation. Moreover, powerful pre-trained forecasting models [timesfm, timer, moirai] can be optionally utilized for better representation projection. Additionally, $w$ represents the top-$K$ similarity scores estimated by the similarity operator $\operatorname{Sim}(\cdot,\cdot)$, which is instantiated using cosine similarity due to its computational efficiency.
Once the relevant time series are retrieved, they serve as references for revising the forecasting results. Instead of modifying the architecture of the backbone forecasting models to incorporate these references, we adopt a practical assumption: similar instances tend to exhibit similar future trends. Therefore, the retrieved references themselves can serve as effective estimations for the target series. This design ensures that the framework remains entirely independent of the backbone models, making it applicable to any forecasting model. In practice, the similarity scores are treated as importance indicators, and the global revised results $y_{global}$ are generated through a weighted sum operation, formulated as follows:
$p$ $=\operatorname{Softmax}(w),$ (4)
$y_{global}$ $=\operatorname{WeightedSum}(p,Y_{re}),$
where the $\operatorname{Softmax(\cdot)}$ function ensures that $\sum_{i=1}^{K}p_{i}=1$, assigning higher weights to retrieved instances that are more similar to the input series.
3.5 Optimization Target
By combining the components described above, the PIR framework produces the final forecasting results through post-forecasting identification and revision using a residual approach [resnet]:
$y_{pred}$ $=\bar{y}+\alpha y_{local}+\beta y_{global},$ (5)
$\alpha$ $=\sigma(\operatorname{Linear}(\delta)),$
$\beta$ $=\sigma(\operatorname{MLP}(\delta,w)).$
Here, $\alpha$ and $\beta$ are learned weights corresponding to the local and global revision components, respectively, and $\sigma(\cdot)$ denotes the Sigmoid activation function. A linear transformation is employed to estimate $\alpha$, with its weight and bias initialized to vectors of ones and zeros. This design ensures ensuring a positive correlation that higher uncertainty estimates lead to larger values of $\alpha$, thus placing greater emphasis on the local revision. In contrast, $\beta$ is generated through a multi-layer perceptron (MLP), which takes both the estimated uncertainty and the retrieval similarities as input. This allows the model to dynamically adjust the influence of the global revision based on the confidence of the prediction and the relevance of the retrieved historical series. The overall optimization objective is to minimize the MSE between the revised forecasts and the ground truth, formulated as:
$\mathcal{L}_{pr}=\frac{1}{N}\sum_{1}^{N}\|y_{pred}-y\|_{2}^{2}.$ (6)
Finally, let $\lambda$ denote the weight hyperparameters for the auxiliary constraint defined in Section 3.2, the overall optimization objective for the PIR framework is then formulated as a multitask learning problem:
$\mathcal{L}=\mathcal{L}_{pr}+\lambda\mathcal{L}_{ue}.$ (7)
4 Experiments
In this section, we conduct extensive experiments on widely used benchmark datasets, comparing our proposed PIR framework with mainstream forecasting approaches under both long-term and short-term forecasting settings, to demonstrate its effectiveness.
Table 1: Forecasting performance under long-term and short-term settings. The results are averaged from all the target series lengths, and the full results are provided in the Appendix. The bold values indicate better performance.
Methods PatchTST + PIR Imp.(%) SparseTSF + PIR Imp.(%) iTrans. + PIR Imp.(%) TimeMixer + PIR Imp.(%)
ETTh1 MSE 0.466 0.437 6.22 0.444 0.433 2.48 0.451 0.432 4.21 0.445 0.429 3.60
MAE 0.452 0.439 2.88 0.431 0.429 0.46 0.445 0.437 1.80 0.436 0.431 1.15
ETTh2 MSE 0.384 0.375 2.34 0.384 0.373 2.86 0.383 0.377 1.57 0.380 0.378 0.53
MAE 0.405 0.400 1.23 0.403 0.398 1.24 0.406 0.403 0.74 0.405 0.403 0.49
ETTm1 MSE 0.397 0.383 3.53 0.416 0.378 9.13 0.407 0.383 5.90 0.381 0.377 1.05
MAE 0.408 0.397 2.70 0.407 0.390 4.18 0.412 0.397 3.64 0.396 0.393 0.76
ETTm2 MSE 0.281 0.283 -0.71 0.287 0.281 2.09 0.291 0.288 1.03 0.276 0.274 0.72
MAE 0.329 0.330 -0.30 0.329 0.328 0.30 0.334 0.334 0.00 0.322 0.323 -0.31
Electricity MSE 0.215 0.200 6.98 0.224 0.196 12.50 0.179 0.175 2.23 0.185 0.181 2.16
MAE 0.303 0.279 7.92 0.297 0.275 7.41 0.269 0.265 1.49 0.275 0.270 1.82
Solar MSE 0.269 0.244 9.29 0.385 0.275 28.57 0.236 0.231 2.12 0.231 0.223 3.46
MAE 0.307 0.287 6.51 0.370 0.296 20.00 0.263 0.260 1.14 0.270 0.267 1.11
Weather MSE 0.259 0.254 1.93 0.276 0.261 5.43 0.260 0.255 1.92 0.245 0.244 0.41
MAE 0.281 0.277 1.42 0.294 0.282 4.08 0.280 0.277 1.07 0.274 0.274 0.00
Traffic MSE 0.482 0.459 4.77 0.637 0.477 25.12 0.425 0.420 1.18 0.519 0.492 5.20
MAE 0.308 0.299 2.92 0.379 0.314 17.15 0.283 0.280 1.06 0.307 0.291 5.21
PEMS03 MSE 0.158 0.120 24.05 0.351 0.154 56.13 0.115 0.107 6.96 0.089 0.089 0.00
MAE 0.265 0.231 12.83 0.400 0.261 34.75 0.225 0.216 4.00 0.200 0.200 0.00
PEMS04 MSE 0.206 0.140 32.04 0.370 0.171 53.78 0.108 0.097 10.19 0.083 0.081 2.41
MAE 0.305 0.251 17.70 0.419 0.278 33.65 0.220 0.206 6.36 0.191 0.190 0.52
PEMS07 MSE 0.165 0.115 30.30 0.352 0.149 57.67 0.095 0.089 6.32 0.087 0.083 4.60
MAE 0.268 0.223 16.79 0.404 0.249 38.37 0.198 0.189 4.55 0.191 0.187 2.09
PEMS08 MSE 0.186 0.147 20.97 0.362 0.180 50.28 0.147 0.135 8.16 0.130 0.128 1.54
MAE 0.284 0.251 11.62 0.413 0.279 32.45 0.245 0.237 3.27 0.238 0.232 2.52
4.1 Experimental Setup
Datasets. For the long-term forecasting task, we conduct experiments on a widely recognized benchmark dataset that includes eight real-world datasets spanning diverse domains [wu2021autoformer, lstnet]. Additionally, we incorporate the PEMS dataset, which contains four subsets, for the short-term forecasting task [scinet]. The exogenous information used in these datasets are the available timestamps. We also conduct experiments on datasets with additional textual descriptions in the Appendix. Following standard experimental protocols, we split each dataset into training, validation, and testing sets in chronological order. The split ratios are set to 6:2:2 for the ETT dataset and 7:1:2 for the other datasets. Detailed information about the datasets is available in the Appendix.
Backbone Models. PIR is a model-agnostic plugin that can be seamlessly integrated with any time series forecasting model. To demonstrate the effectiveness of the framework, we select four mainstream forecasting models based on diverse architectures, encompassing both channel-dependent and channel-independent assumptions, as backbones. These models are evaluated under both long-term and short-term forecasting settings: PatchTST [patchtst], SparseTSF [sparsetsf], iTransformer [itransformer] and TimeMixer [timemixer]. We implement these models following their official code.
Experiments Details. We employ the ADAM optimizer as the default optimization algorithm across all experiments and evaluate performance using two metrics: mean squared error (MSE) and mean absolute error (MAE). For the PIR framework, the retrieval number $K$ is tuned from the set $\{10,20,50\}$, and the weight hyperparameter $\lambda$ is fixed at 1. All experiments are conducted on a single NVIDIA RTX 4090 GPU.
4.2 Main Results
Following the standard evaluation protocol [Informer, yu2024revitalizing], we set the input series length $L_{in}=96$ across all datasets. For a unified comparison, the target series length $L_{out}$ is set to $\{12,24,36,48\}$ for the PEMS dataset and $\{96,192,336,720\}$ for the remaining datasets. The forecasting results for both long-term and short-term settings, along with the corresponding relative performance improvements, are summarized in Table 1.
As shown in the table, the proposed PIR framework consistently improves the performance of state-of-the-art forecasting models across most scenarios in both long-term and short-term forecasting settings. This improvement is primarily due to PIR’s capability to identify potential forecasting failures and effectively leverage contextual information for revision. Specifically, across all 48 experimental settings, PIR achieves an average MSE reduction of 8.99% for PatchTST. Similar trends are observed for other backbone models, with reductions of 25.87% for SparseTSF, 3.47% for iTransformer, and 2.34% for TimeMixer. These results underscore the generalizability of the PIR framework, demonstrating its seamless integration with diverse forecasting models, regardless of their architecture or inductive biases, to deliver enhanced predictive performance.
Additionally, we observe that the relative performance improvements for channel-dependent approaches are smaller compared to those for channel-independent models. This is likely because channel-dependent methods already incorporate covariate information, resulting in stronger baseline performance and leaving less room for improvement. Nevertheless, by leveraging the forecasted covariates, future exogenous variables, and historical time series as contextual information, the PIR framework still manages to enhance the performance of these models. Notably, although the local revision component of PIR shares a structural resemblance with iTransformer, it continues to yield substantial relative improvements. These findings further support our hypothesis that instance-level variations contribute to forecasting failures in specific cases. Moreover, the comparison also highlights the importance of post-forecasting failure identification and revision in generating more reliable and robust forecasting results.
4.3 Qualitative Analysis
Figure 3: Comparison of the groundtruth forecasting error and the estimation of PIR on Solar and Traffic dataset. The backbone is SparseTSF, and the target length is set to 336.
Figure 4: Illustration of the per-instance MSE distribution differences for four mainstream forecasting approaches, both with and without the enhancement of PIR. The MSE values corresponding to the peak density are also highlighted. The comparisons are performed on the ETTh1 and PEMS07 datasets, with the target length set to 96 and 48.
To better illustrate how the PIR framework works to help enhance the forecasting performance, we provide a qualitative analysis of the identification and revision process in this section.
We begin by evaluating the reliability of the failure identification component through a comparison between the actual forecasting error of the backbone models and PIR’s estimated error, $\delta$, as defined in Equation 1. Specifically, we visualize both metrics over a segment of the test set from the Solar and Traffic datasets, using SparseTSF as the backbone model, in Figure 3. The results demonstrate that the PIR framework accurately estimates the forecasting error of the backbone’s intermediate predictions. Notably, the estimated and actual error curves exhibit consistent patterns in terms of peaks and troughs, which validates PIR’s capability to assess the quality of individual forecasting results. This strong alignment underscores the effectiveness of PIR’s uncertainty estimation mechanism and its ability to reliably identify potential forecasting failures, laying a solid foundation for the subsequent revision stage.
For the revision component, although the quantitative results reported in Table 1 already demonstrate its effectiveness in improving overall forecasting performance, we further investigate how the PIR framework mitigates the effects of instance-level variance. Specifically, we evaluate the per-instance MSE of the baseline backbone models on the ETTh1 and PEMS07 datasets. As illustrated in Figure 4, we compare the MSE distribution with and without PIR enhancement, measured using kernel density estimation, and highlight the MSE corresponding to the peak density for each case. The figure reveals that models enhanced by PIR yield more reliable forecasting results, as their error distributions exhibit higher density at lower MSE values. Additionally, the MSE corresponding to the peak density is significantly reduced. Additionally, PIR framework also effectively addresses forecasting failures, substantially reducing errors for instances poorly modeled by the backbone models. This improvement is particularly evident in the PEMS07 dataset, where the tail of the error distribution curve shifts significantly toward smaller MSE values. For example, the maximum prediction error of SparseTSF on ETTh1 decreases from 2.85 to 0.81 when enhanced with PIR.
In summary, the qualitative analysis strongly aligns with our expectations. The PIR framework can effectively identify forecasting failures and enhance overall performance by revising the intermediate results in a model-agnostic manner.
4.4 Complexity Analysis
In this section, we analyze the computational overhead of the proposed PIR framework from both theoretical and empirical perspectives. Theoretically, the series-wise cosine similarity function used in the retrieval stage has a time complexity of $O(NML_{in})$, where $N$ denotes the number of channels, $M$ is the total number of historical series per channel, and $L_{in}$ represents the input sequence length. The subsequent local revision process involves channel-wise attention operations, resulting in a complexity of $O(N^{2})$. To further evaluate the practical efficiency of PIR, we report the average inference time on both a small-scale dataset (ETTh1) and a large-scale dataset (Traffic). We compare two retrieval strategies: brute-force cosine similarity and approximate retrieval using Locality-Sensitive Hashing (LSH). The results are summarized in Table 2.
Table 2: Inference time comparison under different settings.
ETTh1(Cos) ETTh1(LSH) Traffic(Cos) Traffic(LSH)
Backone(s) 0.164 0.164 0.424 0.424
[…]
MSE Improvement 0.014 0.009 0.025 0.025
The results indicate that the retrieval stage introduces negligible additional latency on both datasets, thanks to the GPU-parallelizable nature of cosine similarity. For even larger datasets, the total computational cost can be further reduced by applying sampling strategies (e.g., stride sampling) or dimensionality reduction techniques to limit the search space. In contrast, the LSH-based retrieval implemented with the faiss library yields significantly higher inference time without performance gains, indicating that brute-force cosine similarity is both more efficient and effective in our current implementation.
5 Conclusion
In this paper, we first investigated the challenge of instance-level variance in time series forecasting, which often results in unreliable predictions for certain cases. To address this, we proposed the PIR framework, a model-agnostic solution that enhances the accuracy through post-forecasting identification and revision of potentially biased predictions. The framework identifies the forecasting failures by estimating their error on a per-instance basis, and leverages contextual information to revise forecasts from both local and global perspectives. Extensive experiments across various benchmarks and forecasting models demonstrated that PIR consistently improves performance, highlighting its effectiveness and versatility as a plug-in component for a wide range of forecasting architectures. We wish our work could raise new research directions for the forecasting task.
6 Acknowledgement
This work was supported by the National Natural Science Foundation of China (No. U23A20319, No. 62502486). We furthermore thanked the anonymous reviewers for their constructive comments.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We provide clear descriptions of our contributions and scope in the abstract and introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We provide discussions on the limitations of our work in Section  F.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: We don’t claim any theorems.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide the implementation and running scripts that ensure the reproducibility in https://anonymous.4open.science/r/PIR-70BF/.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide the links to the data we used in the paper in the Appendix, and our code in https://github.com/icantnamemyself/PIR.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We specify all the training and test details in Section 4.1.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: Due to layout factors, we don’t provide the experiment statistical significance.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: As discussed in Section 4.1, all experiments can be reproduced using on a single NVIDIA RTX 4090 GPU.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We review and follow the NeurIPS Code of Ethics to conduct our reseach.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: As discussed in Section 5, our work could enhance the performance of time series forecasting, which is a fundamental task in everyday life. Besides, we wish our work could raise new research directions for the forecasting task.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our work poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The code and data we used are publicly available, and we cite the original papers properly.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We release our code for the proposed method.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We don’t involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We don’t involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: We only use LLMs to modify the grammar and refine the text.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Dataset Descriptions
In this paper, we leverage a diverse set of forecasting datasets covering various domains to evaluate the effectiveness of the proposed PIR framework under both long-term and short-term forecasting settings. We also include datasets with additional textual information [timemmd] to validate the framework’s generalizability. The brief descriptions and characteristics are presented as follows:
  • ETT [^2] https://github.com/zhouhaoyi/ETDataset: The dataset records oil temperature and load metrics from electricity transformers, tracked between July 2016 and July 2018. It is subdivided into four mini-datasets, with data sampled either hourly or every 15 minutes.
  • Electricity [^3] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014: The dataset captures the hourly electricity consumption in kWh of 321 clients, monitored from July 2016 to July 2019.
  • Solar [^4] http://www.nrel.gov/grid/solar-power-data.html: The dataset records the solar power production in the year of 2006, which is sampled every 10 minutes from 137 PV plants in Alabama State.
  • Weather [^5] https://www.bgc-jena.mpg.de/wetter/: The dataset records the 21 weather indicators, including air temperature and humidity every 10 minutes from the Weather Station of the Max Planck Biogeochemistry Institute in 2020.
  • Traffic [^6] http://pems.dot.ca.gov: The dataset provides the hourly traffic volume data describing the road occupancy rates of San Francisco freeways, recorded by 862 sensors.
  • PEMS [^7] https://github.com/guoshnBJTU/ASTGNN/tree/main/data: The dataset is a series of traffic flow dataset with four subsets, including PEMS03, PEMS04, PEMS07, and PEMS08. The traffic information is recorded at a rate of every 5 minutes by multiple sensors.
  • Energy and Health [^8] https://github.com/AdityaLab/MM-TSFlib: These two datasets are subsets of Time-MMD [timemmd], a multimodal time series dataset that ensures fine-grained alignment between textual and numerical modalities. The datasets are collected weekly, spanning from 1996 and 1997 up to May 2024, respectively.
Table 3: The overview of each dataset used in the experiments.
Dataset Variables Frequency Length Scope
ETTh1&ETTh2 7 1 Hour 17420 Energy
ETTm1&ETTm2 7 15 Minutes 69680 Energy
Electricity 321 1 Hour 26304 Energy
Solar 137 10 Minutes 52560 Nature
Weather 21 10 Minutes 52696 Nature
Traffic 862 1 Hour 17544 Transportation
PEMS03 358 5 Minutes 26208 Transportation
PEMS04 307 5 Minutes 16992 Transportation
PEMS07 883 5 Minutes 28224 Transportation
PEMS08 170 5 Minutes 17856 Transportation
Energy 9 1 Week 1479 Energy
Health 11 1 Week 1389 Health
Appendix B Full Experimental Results
In this section, we present the complete long-term and short-term forecasting results in Table 4. These results demonstrate that the proposed PIR framework functions as a model-agnostic plugin, consistently improving the forecasting performance of various backbone models across diverse datasets and settings.
Table 4: Full experimental results of long-term and short-term forecasting. The target length $L_{out}$ is chosen as {12,24,36,48} for the PEMS dataset and {96,192,336,720} for the others. The bold values indicate better performance.
Methods PatchTST + PIR SparseTSF + PIR iTransformer + PIR TimeMixer + PIR
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1
96 0.410 0.416 0.375 0.400 0.399 0.401 0.375 0.392 0.385 0.403 0.376 0.402 0.384 0.399 0.370 0.397
192 0.458 0.443 0.422 0.427 0.438 0.423 0.420 0.416 0.440 0.434 0.424 0.429 0.432 0.425 0.422 0.421
336 0.498 0.464 0.467 0.451 0.478 0.443 0.465 0.440 0.486 0.457 0.465 0.450 0.486 0.449 0.460 0.442
720 0.498 0.486 0.484 0.477 0.461 0.457 0.473 0.466 0.494 0.485 0.461 0.466 0.476 0.470 0.464 0.462
ETTh2
96 0.299 0.347 0.291 0.340 0.304 0.347 0.289 0.338 0.302 0.350 0.298 0.346 0.295 0.343 0.291 0.343
192 0.384 0.398 0.371 0.391 0.386 0.396 0.373 0.392 0.382 0.399 0.374 0.396 0.378 0.398 0.367 0.395
336 0.424 0.432 0.418 0.429 0.424 0.429 0.418 0.426 0.423 0.433 0.415 0.429 0.421 0.436 0.422 0.431
720 0.427 0.444 0.421 0.441 0.421 0.438 0.412 0.435 0.424 0.443 0.420 0.440 0.427 0.441 0.430 0.444
ETTm1
96 0.336 0.375 0.317 0.354 0.357 0.375 0.311 0.350 0.342 0.377 0.315 0.356 0.316 0.355 0.309 0.351
192 0.376 0.394 0.365 0.383 0.394 0.392 0.355 0.374 0.381 0.395 0.358 0.380 0.364 0.383 0.359 0.381
336 0.409 0.416 0.394 0.404 0.426 0.414 0.389 0.396 0.420 0.420 0.395 0.406 0.389 0.403 0.387 0.402
720 0.465 0.445 0.457 0.445 0.486 0.447 0.457 0.438 0.486 0.456 0.465 0.447 0.455 0.441 0.451 0.440
ETTm2
96 0.177 0.263 0.175 0.259 0.185 0.267 0.174 0.258 0.184 0.267 0.179 0.263 0.175 0.257 0.170 0.254
192 0.242 0.305 0.241 0.306 0.248 0.306 0.239 0.301 0.254 0.313 0.246 0.307 0.239 0.299 0.239 0.300
336 0.304 0.345 0.304 0.347 0.308 0.343 0.304 0.345 0.312 0.350 0.311 0.351 0.298 0.339 0.296 0.340
720 0.401 0.401 0.410 0.409 0.408 0.398 0.408 0.407 0.412 0.407 0.417 0.415 0.393 0.394 0.389 0.397
Electricity
96 0.193 0.284 0.180 0.253 0.210 0.280 0.174 0.250 0.148 0.240 0.145 0.237 0.156 0.248 0.151 0.244
192 0.198 0.289 0.184 0.262 0.206 0.281 0.180 0.259 0.163 0.254 0.161 0.251 0.170 0.261 0.165 0.258
336 0.214 0.304 0.198 0.278 0.219 0.296 0.195 0.276 0.177 0.270 0.175 0.268 0.187 0.277 0.186 0.275
720 0.255 0.336 0.239 0.322 0.260 0.329 0.233 0.314 0.227 0.311 0.219 0.303 0.227 0.312 0.221 0.307
Solar
96 0.233 0.287 0.210 0.263 0.336 0.351 0.231 0.270 0.205 0.238 0.196 0.235 0.198 0.261 0.195 0.248
192 0.266 0.307 0.241 0.286 0.376 0.371 0.272 0.292 0.237 0.262 0.235 0.257 0.241 0.274 0.238 0.275
336 0.291 0.317 0.261 0.297 0.415 0.384 0.301 0.313 0.251 0.275 0.248 0.275 0.253 0.274 0.235 0.273
720 0.286 0.316 0.263 0.300 0.413 0.374 0.297 0.310 0.250 0.276 0.246 0.274 0.231 0.271 0.225 0.271
Weather
96 0.179 0.220 0.168 0.208 0.201 0.240 0.175 0.218 0.174 0.214 0.170 0.211 0.163 0.210 0.161 0.208
192 0.225 0.259 0.216 0.253 0.242 0.273 0.220 0.256 0.222 0.255 0.217 0.252 0.208 0.252 0.206 0.250
336 0.279 0.298 0.276 0.297 0.294 0.308 0.284 0.302 0.281 0.299 0.277 0.298 0.266 0.291 0.263 0.291
720 0.354 0.347 0.355 0.350 0.366 0.355 0.364 0.353 0.361 0.352 0.356 0.350 0.341 0.343 0.344 0.347
Traffic
96 0.459 0.298 0.428 0.288 0.664 0.395 0.454 0.306 0.393 0.268 0.390 0.266 0.489 0.296 0.453 0.275
192 0.468 0.301 0.450 0.294 0.611 0.366 0.456 0.302 0.415 0.277 0.415 0.278 0.495 0.299 0.473 0.285
336 0.483 0.307 0.466 0.299 0.619 0.367 0.480 0.313 0.429 0.284 0.426 0.283 0.533 0.311 0.503 0.294
720 0.518 0.326 0.493 0.315 0.655 0.387 0.516 0.333 0.461 0.301 0.447 0.299 0.557 0.322 0.539 0.309
PEMS03
12 0.085 0.196 0.074 0.183 0.145 0.258 0.081 0.191 0.069 0.174 0.067 0.172 0.063 0.168 0.064 0.170
24 0.135 0.249 0.105 0.217 0.267 0.352 0.126 0.238 0.098 0.209 0.092 0.202 0.084 0.195 0.084 0.196
36 0.180 0.287 0.134 0.246 0.416 0.449 0.178 0.285 0.130 0.243 0.119 0.231 0.099 0.213 0.099 0.212
48 0.231 0.326 0.165 0.276 0.574 0.541 0.229 0.328 0.164 0.275 0.149 0.260 0.110 0.224 0.110 0.223
PEMS04
12 0.106 0.218 0.089 0.198 0.158 0.274 0.096 0.207 0.081 0.189 0.077 0.182 0.070 0.174 0.069 0.173
24 0.170 0.279 0.122 0.235 0.286 0.374 0.143 0.256 0.099 0.211 0.090 0.199 0.079 0.186 0.078 0.185
36 0.239 0.337 0.156 0.268 0.436 0.468 0.194 0.301 0.119 0.233 0.105 0.216 0.086 0.195 0.085 0.193
48 0.310 0.386 0.191 0.301 0.601 0.561 0.250 0.347 0.134 0.248 0.116 0.227 0.097 0.210 0.094 0.207
PEMS07
12 0.080 0.190 0.070 0.175 0.131 0.248 0.075 0.180 0.066 0.160 0.065 0.161 0.059 0.157 0.057 0.155
24 0.134 0.246 0.101 0.210 0.262 0.354 0.120 0.228 0.088 0.191 0.083 0.183 0.077 0.181 0.076 0.179
36 0.193 0.295 0.130 0.240 0.423 0.459 0.167 0.272 0.105 0.211 0.097 0.199 0.095 0.201 0.094 0.200
48 0.253 0.339 0.157 0.266 0.593 0.556 0.233 0.314 0.121 0.228 0.110 0.213 0.115 0.224 0.106 0.214
PEMS08
12 0.097 0.208 0.088 0.194 0.150 0.266 0.094 0.201 0.081 0.185 0.079 0.184 0.079 0.184 0.078 0.182
24 0.153 0.259 0.126 0.233 0.274 0.365 0.148 0.254 0.123 0.227 0.115 0.221 0.113 0.224 0.111 0.218
36 0.217 0.313 0.168 0.272 0.425 0.462 0.205 0.305 0.167 0.264 0.153 0.256 0.142 0.250 0.141 0.246
48 0.278 0.356 0.207 0.304 0.598 0.559 0.272 0.357 0.217 0.305 0.194 0.288 0.184 0.292 0.180 0.283
Appendix C Generalizability Investigation
In this section, we conduct comparative experiments on the Energy and Health datasets, which include additional aligned textual descriptions as exogenous information. These experiments aim to assess the generalizability of the proposed PIR framework in handling more complex contextual information. The results, presented in Table 5, show that PIR consistently improves performance in most cases, thereby validating its effectiveness in multimodal forecasting scenarios.
Table 5: Comparison experiments on the Energy and Health datasets. The input series length $L_{in}$ is set to 24, and the target length $L_{out}$ is chosen as {12,24,36,48}. The bold values indicate better performance.
Methods PatchTST + PIR SparseTSF + PIR iTransformer + PIR TimeMixer + PIR
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Energy
12 0.194 0.334 0.130 0.255 0.140 0.273 0.128 0.255 0.122 0.245 0.117 0.242 0.149 0.286 0.135 0.263
24 0.299 0.418 0.248 0.364 0.252 0.375 0.241 0.362 0.242 0.365 0.232 0.355 0.226 0.348 0.222 0.346
36 0.382 0.475 0.337 0.432 0.335 0.436 0.321 0.419 0.327 0.426 0.309 0.414 0.319 0.422 0.314 0.420
48 0.463 0.527 0.425 0.496 0.418 0.493 0.416 0.493 0.420 0.493 0.406 0.485 0.408 0.487 0.393 0.477
Health
12 12.493 1.960 9.901 1.612 11.333 1.802 10.315 1.641 8.523 1.469 8.258 1.461 8.947 1.493 8.642 1.489
24 15.252 2.204 13.486 1.924 14.115 2.077 13.837 1.975 12.159 1.801 11.698 1.748 12.010 1.789 11.790 1.765
36 13.804 2.130 12.032 1.893 12.856 2.052 12.876 1.969 11.662 1.854 11.093 1.793 11.385 1.801 11.029 1.792
48 12.515 2.062 12.294 1.965 13.167 2.015 11.784 2.001 10.910 1.817 10.332 1.781 11.227 1.879 10.328 1.784
Appendix D Ablation Study
In this section, we investigate the impact of the auxiliary constraint and structural designs on the overall performance of the PIR framework through ablation studies. The backbone models selected for this analysis are PatchTST and iTransformer, representing the channel-independent and channel-dependent categories, respectively.
[…]
Table 6: MAE comparison between different variants of the PIR framework. The Bold values indicate the best performance.
Backbone PatchTST iTransformer
Variants w/o PIR w/o Local w/o Global w/ PIR w/o PIR w/o Local w/o Global w/ PIR
Electricity
96 0.284 0.257 0.270 0.253 0.240 0.238 0.241 0.237
192 0.289 0.266 0.277 0.262 0.254 0.253 0.256 0.251
336 0.304 0.282 0.292 0.278 0.270 0.268 0.272 0.268
720 0.336 0.326 0.325 0.322 0.311 0.305 0.304 0.303
PEMS03
12 0.196 0.201 0.192 0.183 0.174 0.175 0.174 0.172
24 0.249 0.243 0.228 0.217 0.209 0.204 0.207 0.202
36 0.287 0.281 0.262 0.246 0.243 0.234 0.239 0.231
48 0.326 0.314 0.298 0.276 0.275 0.262 0.272 0.260
To evaluate the impact of the auxiliary constraint, we compare forecasting performance across various values of the weight hyperparameter $\lambda$, ranging from 0 to 1. Additionally, we report the corresponding uncertainty estimation error $\mathcal{L}_{ue}$, as defined in Equation 1, in Figure 5. The results suggest that the auxiliary constraint enables the PIR framework to more accurately estimate the uncertainty of intermediate forecasts by predicting their associated errors, thereby facilitating the revision weight learning, leading to better accuracy. Furthermore, the R-squared scores between MSE and $\mathcal{L}_{ue}$ are 0.9067 and 0.7500 when PatchTST and iTransformer are used as the backbone models, respectively. These high correlations strongly support the feasibility of our approach, validating the use of forecasting error as a proxy for uncertainty in time series forecasting tasks.
On the other hand, we present the MAE comparison between different variants of the PIR framework on the Electricity and PEMS03 datasets in Table 6, providing an intuitive understanding of how the local and global revision components contribute to the performance. It can be inferred from the results that both components contribute to better performance compared to the baseline model under most cases, and the combination of them consistently leads to the best accuracy, validating our proposal that utilizing contextual information to revise forecasting results from both local and global perspectives can well alleviate the impact of the instance-level variance.
To further examine where the performance gains of the PIR framework originate, we conduct an ablation study comparing PIR with two variants: (1) a deepened iTransformer, which increases model depth to match the additional layers introduced by PIR’s local revision component, and (2) PIR trained from scratch, which removes the requirement of a pretrained forecasting backbone. The results on the Electricity and Weather datasets are presented in Table 7. The results show that PIR consistently outperforms both variants. Benefiting from both global and local revision components, PIR effectively utilizes retrieved similar historical series as well as valuable contextual insights from covariates and exogenous variables. This allows PIR to outperform simply enlarging model capacity. Furthermore, training PIR from scratch leads to degraded performance, possibly because the randomly initialized backbone produces unstable forecasts in early training stages, which negatively affects the optimization of the failure identification module and weakens its ability to estimate forecasting error.
Table 7: Investigation into the sources of performance improvement. The Bold values indicate the best performance.
Variants iTransformer w/ PIR w/ Deepen w/ PIR scratch
Metric MSE MAE MSE MAE MSE MAE MSE MAE
Electricity
96 0.148 0.240 0.145 0.237 0.147 0.239 0.164 0.256
192 0.163 0.254 0.161 0.251 0.163 0.255 0.178 0.270
336 0.177 0.270 0.175 0.268 0.177 0.271 0.196 0.284
720 0.227 0.311 0.219 0.303 0.217 0.304 0.256 0.334
Weather
96 0.174 0.214 0.170 0.211 0.177 0.218 0.217 0.249
192 0.222 0.255 0.217 0.252 0.224 0.257 0.252 0.270
336 0.281 0.299 0.277 0.298 0.279 0.296 0.300 0.315
720 0.361 0.352 0.356 0.350 0.358 0.350 0.393 0.371
Appendix E Forecasting showcases
In this section, we provide intuitive forecasting showcases to illustrate the impact of the revision process on forecasting performance, and the effect of global revision in cases with and without similar historical instances.
Figure 6: Illustration of forecasting showcases of different variants of PIR using PatchTST as the backbone model.
In Figure 6, we examine the impact of the revision process on forecasting results using PatchTST as the backbone model across three datasets of varying scales. For the local revision component, it struggles to capture scale changes using only cross-channel dependencies and exogenous information on the ETTh1 dataset, but successfully corrects trend deviations to better align with the ground truth on the Solar dataset. Meanwhile, the global revision component improves future scale predictions by retrieving similar historical series on the ETTh1 dataset but has minimal impact on the Solar dataset. Furthermore, both local and global revisions contribute to improved forecasting accuracy on the PEMS04 dataset, with their combination (i.e., the PIR framework) achieving superior performance. These findings indicate that local and global revisions each have their strengths and limitations, underscoring the importance of constructing a unified framework that effectively integrates both approaches.
Figure 7: Illustration of forecasting showcases on various forecasting instances with and without similar historical time series.
Additionally, we evaluate the performance of global revision with and without similar historical instances. In Figure 7, we present the retrieval results ( $y_{global}$, as defined in Equation 4 of our paper) as the forecasting outputs, along with the corresponding average top-k similarities and learned weight $\beta$. The results indicate that retrieval similarities vary significantly across instances. For instances with highly similar historical series, the retrieval process effectively captures future evolutions, whereas for those without close historical analogs, the retrieval results only approximate future trends and may even be inaccurate. This observation aligns with our expectations, as the retrieval mechanism is designed to leverage similar past series as references for forecasting. Furthermore, the PIR framework adaptively assigns higher weights to instances with stronger retrieval similarities, dynamically adjusting the influence of retrieval results on the final forecast.
Appendix F Limitations
Though PIR demonstrates promising performance on benchmark datasets, there are still several limitations within this framework. Firstly, the instance-level variances exist in both input series and target series, while the framework currently addresses only the former challenge. Identifying and addressing instance-level variances in the target series caused by data quality issues—such as noise, outliers, and missing data—holds significant potential for improving both the training and evaluation processes. Furthermore, relatively simple network architectures are utilized in the failure identification and local revision components. Exploring ways to integrate advanced inductive biases into these components to enhance their performance remains an important direction for future research.